OpenStudy (anonymous):

Pls help me...Circles A and B have radii 5 and 7, respectively, and the distance between their centers A and B is 6. If the two circles intersect at points M and N, find the length of the common chord MN.

OpenStudy (anonymous):

For ALL geometry questions you MUST make a drawing. Have you done that already?

OpenStudy (anonymous):



OpenStudy (anonymous):

Something like this, preferable on paper with a pencil

OpenStudy (anonymous):

yup im done...

OpenStudy (anonymous):

Hm wait a moment, I will make a drawing on paper, this doesnt help me so much

OpenStudy (anonymous):

is x = 1?

OpenStudy (anonymous):

i mean h =2 sqrt 6...

OpenStudy (anonymous):

so MN = 4 sqrt 6?

OpenStudy (anonymous):

I dont know yet. But I would prefer if you posted your way of solving it instead of just posting numbers

OpenStudy (anonymous):

Yes answer is 4sqrt6

OpenStudy (anonymous):

equate area of triangle 1/2*h*6=area by herons formula

OpenStudy (anonymous):

ok thank you so much...i must only use the pythagorean theorem..right?

OpenStudy (anonymous):

This is possible since common chord is always perpendicular to the line joining centres of a circle and is bisected by it.

OpenStudy (anonymous):

yep...but at first it seems impossible since the distance of the 2 centers is 6 yet the radii are 5 and 7...

OpenStudy (anonymous):

how to apply heron's formula here MayankD?

OpenStudy (anonymous):

Yea the figure is somewhat like 

OpenStudy (anonymous):

yep dts ryt..

OpenStudy (anonymous):

Area of a triangle=sqrt(s*(s-a)*(s-b)*(s-c)) where s is semi perimeter of triangle and a b and c are the sides of the traingle

OpenStudy (anonymous):

Yes I have \[4\sqrt{6}\] as well.

OpenStudy (anonymous):

Please help me... Suppose f(t)=A*2^(0.02t) represents the bacteria in a culture t minutes from the start of an experiment. Give an expression for the number of minutes needed for the bacteria to double.HW BWT DS??

OpenStudy (anonymous):

F(t)=A*2^(0.02(t+50))

OpenStudy (anonymous):

d answer here is 50ln2 but i dont know the solution

OpenStudy (anonymous):

What is the exact question?